OpenStudy (anonymous):

Use L'hospital rule where applicable lim √x(e^x/2) x->∞

zepdrix (zepdrix):

Everything under the root, or just x?

OpenStudy (anonymous):

just x

OpenStudy (anonymous):

root x time e^(-x/2)

zepdrix (zepdrix):

As \(\Large x\to\infty\), our function is approaching the indeterminate form \(\Large \infty\cdot0\) right?

OpenStudy (anonymous):

right

zepdrix (zepdrix):

Using rules of exponents, \[\Large \lim_{x\to\infty}\sqrt x \;e^{-x/2}\quad=\quad \lim_{x\to\infty}\frac{\sqrt x}{e^{x/2}}\]

zepdrix (zepdrix):

Now it's giving us the indeterminate form \(\Large \dfrac{0}{0}\). Good! Now it's in a nice form where we can apply L'Hop, yes?

zepdrix (zepdrix):

Sorry I mean infty/infty*

OpenStudy (anonymous):

yeah thats as far as i got on my own. im not 100% confident where to go from there

zepdrix (zepdrix):

So applying L'Hop,\[\Large L'H\quad=\quad \lim_{x\to\infty}\frac{(\sqrt{x})'}{\left(e^{x/2}\right)'}\]What derivatives do we get? :o Remember the derivative of sqrt x?

OpenStudy (anonymous):

yeah its x^1/2

zepdrix (zepdrix):

So that's sqrt x rationalized :D But what is its derivative?

zepdrix (zepdrix):

Power rule, yes?

OpenStudy (anonymous):

1/2 x ^-1/2

zepdrix (zepdrix):

Ok good! We'll write it like this,\[\Large \left(\sqrt x\right)'\quad=\quad \frac{1}{2\sqrt x}\]so it's a bit easier to work with.

zepdrix (zepdrix):

So our derivatives should look like this:\[\Large L'H\quad=\quad \lim_{x\to\infty}\frac{(\sqrt{x})'}{\left(e^{x/2}\right)'}\quad=\quad \lim_{x\to\infty}\frac{\left(\dfrac{1}{2\sqrt x}\right)}{\left(\frac{1}{2}e^{x/2}\right)}\]

OpenStudy (anonymous):

next step would be to multiply by 1/2 correct?

zepdrix (zepdrix):

Multiply the top and bottom by 2? Sure! That might clean things up a tad.

zepdrix (zepdrix):

We can also bring the sqrtx down into the denominator.\[\Large \lim_{x\to\infty}\frac{-1}{\sqrt x\;e^{-x/2}}\]Woops I missed a negative when I took the derivative of the exp part. Fixed it.

zepdrix (zepdrix):

Hmmm that doesn't seem to have helped us much -_- hmmmmm

OpenStudy (anonymous):

what i did was \[1/2\sqrt{x *e^x}\]

zepdrix (zepdrix):

Oh I didn't miss a negative did I? It became positive when we brought it into the denom* blah.

OpenStudy (anonymous):

and then plug in the infinity to get 1/2* infinite * infinite = 1/infinite = 0

OpenStudy (anonymous):

is that correct

zepdrix (zepdrix):

\[\Large \lim_{x\to\infty}\frac{1}{\sqrt {x\;e^{x}}}\]

zepdrix (zepdrix):

I think you have an extra factor of 1/2 in there, but ya looks like you've got it figured out. For some reason I was thinking infty * infty was an indeterminate form :)

zepdrix (zepdrix):

0 is correct, yay good job!

OpenStudy (anonymous):

thanks for guiding me